Let h(x) =f(x)/g(x) where h(x) is polynomial of x. f f(x) = x3-2x2-3x and g(x)=x-3,then find the number of real solutions of the equation h(x)=0 in the interval [-2.0, 2.0]












1
Expert's answer
2021-10-04T16:41:26-0400
"h(x)=\dfrac{x^3-2x^2-3x}{x-3}=\dfrac{x(x^2-2x-3)}{x-3}"

"=\dfrac{x(x+1)(x-3)}{x-3}=x(x+1), x\not=3"

"h(x)=0=>x(x+1)=0, x\not=3"

"x_1=-1, x_2=0"

There are two real solutions of the equation h(x)=0 in the interval [-2.0, 2.0].